a total of $1150 was invested, part of it at 12% and part at 11%. the total yield was $133.75. how much invested at each rate?
Answer by ankor@dixie-net.com(12695)

a total of $1150 was invested, part of it at 12% and part at 11%. the total yield was $133.75. how much invested at each rate?
:
We know the total of the two investments is $1150, so if the 12% amt = x,
Then the 11% amt = (1150-x)
:
Int @ 12% + Int @ 11% = 133.75
:
.12x + .11(1150-x) = 133.75
:
.12x + 126.5 - .11x = 133.75
:
.12x - .11x = 133.75 - 126.5
:
.01x = 7.25
:
x = 7.25/.01
:
x = $725 invested at 12%
:
1150 - 725 = $425 invested at 11%
:
:
Check:
.12(725) + .11(425) =
87 + 46.75 = 133.75